# Efficient and Stealthy Jailbreak Attacks via Adversarial Prompt Distillation from LLMs to SLMs

### v1

@@ Abstract
WARNING: This paper contains unfiltered content generated by LLMs that may be offensive to readers.
Attacks on large language models (LLMs) in jailbreaking scenarios raise many security and ethical issues. Current jailbreak attack methods face problems such as low efficiency, high computational cost, and poor cross-model adaptability and versatility, which make it difficult to cope with the rapid development of LLM and new defense strategies. Our work proposes an Adversarial Prompt Distillation  [^1] Our code is available at: https://github.com/lxgem/Efficient_and_Stealthy_Jailbreak_Attacks_via_Adversarial_Prompt., which combines masked language modeling, reinforcement learning, and dynamic temperature control through a prompt generation and distillation method. It enables small language models (SLMs) to jailbreak attacks on mainstream LLMs. The experimental results verify the superiority of the proposed method in terms of attack success rate and harm, and reflect the resource efficiency and cross-model adaptability. This research explores the feasibility of distilling the jailbreak ability of LLM to SLM, reveals the model’s vulnerability, and provides a new idea for LLM security research.
@@ 1 Introduction
Large language models (LLMs) face serious jailbreak threats that their widespread deployment has exposed. Although various defense strategies are being adopted to ensure the security and reliability of models against jailbreak threats Yi et al. (2024), their vulnerability to adversarial attacks, especially attacks, has become a problem that cannot be ignored. Jailbreak attacks pose significant ethical and security challenges by bypassing security mechanisms and inducing models to generate unexpected or harmful responses, exploiting weaknesses in model design, training data, or input processing Hadi et al. (2023).
Figure 1: The attack text is refined through distillation and mask rewriting, enhancing its stealth and making it more challenging for detection systems to identify malicious intent.
Current jailbreak attack methods are categorized into manually crafted prompts and template-based automated generation. Manually crafted prompts depend on attackers’ or researchers’ domain knowledge to iteratively design inputs that bypass safety mechanisms Shen et al. (2024); Liu et al. (2023c); Russinovich et al. (2024), often using specific sentence structures or semantic manipulation to elicit prohibited content. Although effective in targeted scenarios, these methods are inefficient and difficult to scale due to their reliance on human effort. Conversely, template-based automated generation employs predefined templates and task instructions to create diverse attack promptsLiu et al. (2023b); Schwartz et al. (2025); Li et al. (2024). Recent advancements have incorporated universal template libraries, heuristic selection of high-success-rate prompts, and dynamic optimization through probabilistic sampling. The common drawback of these two types of methods is the inefficiency of generating jailbreak attacks. The high time and space complexity of generative and heuristic methods limits the adaptability of these methods to jailbreak across scenarios.
@@
To overcome these limitations, we introduce Adversarial Prompt Distillation (APD), a multi-stage framework that fundamentally rethinks how to efficiently transfer and optimize jailbreak capabilities. Our approach is based on three key innovations: (1) We build on SLM BERT Devlin et al. (2019) by fine-tuning LoRA Hu et al. (2022) to enable small language models (SLMs) to acquire complex jailbreak capabilities from large language models. This is achieved by combining a masked language modeling objective and a carefully designed KL divergence loss function, which significantly reduces model size and computational requirements while maintaining attack effectiveness.(2) We develop a dynamic temperature control mechanism that automatically adjusts the exploration-exploitation trade-off during prompt generation. This innovation enables the system to maintain an optimal balance between attack stealth (avoiding detection by security filters) and effectiveness (achieving a high success rate). (3) We incorporate a reinforcement learning component that continuously optimizes template selection and combination strategies based on real-time feedback from the target model. This adaptive component enables the system to continuously improve its attack strategy based on changes in defense countermeasures or model behavior. Experimental evaluation confirms the superiority of the APD method in attack success rate and attack efficiency, and achieves the model’s balance of efficiency and low resource usage.
@@
  • We implement the first transfer of generative jailbreaking methods from LLM to SLM, enabling BERT and other SLMs to perform better jailbreaking tasks through knowledge distillation.
  • We propose an efficient multi-stage APD framework that combines masked language modeling, KL divergence loss, and dynamic temperature function to complete knowledge transfer from LLM to SLM efficiently.
  • We verify the superior performance of the APD method, in the AdvBench test, APD is better than the existing methods in attack success rate, harmfulness, and efficiency.
@@ 2 Related Work
2.1 Jailbreak Attacks on LLMs
Recent research has systematically explored the evolving landscape of jailbreak attacks on large language models (LLMs), revealing both sophisticated attack methods and persistent vulnerabilities. The field has progressed from early demonstrations of manual prompt engineering techniques Perez et al. (2022) to advanced automated approaches, including gradient-based adversarial suffix generation Mehrotra et al. (2023) and black-box evolutionary optimization methods Huang et al. (2024). These attacks have proven remarkably transferable across different LLM architectures and providers Deng et al. (2023), highlighting fundamental weaknesses in current alignment approaches. Particularly concerning are emerging attack vectors, such as the gradual erosion of safeguards through multi-turn conversations Greshake et al. (2024), cross-lingual attack strategies that bypass English-centric filters Lin et al. (2024), and the stealthiness embedding of malicious prompts within complex documents or code segments Liu et al. (2024). While defense mechanisms have advanced in parallel-including real-time prompt classification systems Helbling et al. (2024) and adversarially trained alignment techniques Qi et al. (2023), recent evaluations show these protections remain vulnerable to adaptive attackers who continuously refine their methods Wang et al. (2023). This ongoing arms race underscores the need for more robust and fundamental solutions to LLM security challenges.
@@ 2.2 Knowledge Distillation
Knowledge distillation has evolved significantly from its initial formulation using KL divergence for output distribution alignment Hinton et al. (2015) to more sophisticated approaches incorporating intermediate feature matching Romero et al. (2014) and relational knowledge transfer Heo et al. (2019). Recent advances include data-efficient methods like generative adversarial networks for synthetic sample generation Park et al. (2019) and deep inversion techniques for parameter-based reconstruction Lopes et al. (2017), enabling distillation without original training data. The technique has demonstrated versatility across modalities, from cross-modal RGB-depth transfer Yin et al. (2020) to language model compression Hafner et al. (2022), while proving particularly valuable for recommender systems through applications in collaborative filtering Vakili et al. (2024) and graph neural network compression Kang et al. (2020). These developments have expanded the applicability of knowledge distillation while addressing critical challenges in model efficiency and deployment.
@@ 2.3 Reinforcement Learning for LLMs
Recent advances in reinforcement learning for large language models (LLMs) have significantly improved their alignment with human preferences, beginning with the foundational RLHF framework that trains reward models from pairwise human preferences Christiano et al. (2017). This approach was successfully adapted to language models, demonstrating notable improvements in output quality through RL fine-tuning Stiennon et al. (2020). Subsequent refinements, such as Proximal Policy Optimization (PPO), enabled more stable training Ziegler et al. (2019). Recent work has introduced Direct Preference Optimization (DPO) to address efficiency limitations, which bypasses explicit reward modeling to optimize human preferences Rafailov et al. (2023) directly. Meanwhile, other studies have explored self-supervised alignment through Constitutional AI to reduce the need for human annotations Abiri (2024). Scaling challenges have been addressed in large-scale implementations, such as ChatGPT Ouyang et al. (2022). However, persistent issues include reward hacking Ethayarajh et al. (2021) and the need for robust preference learning methods Ethayarajh et al. (2021). The field continues to evolve with techniques like reinforcement learning from AI feedback (RLAIF) that aim to automate alignment while preserving safety Liu et al. (2023a).
@@ 3 Preliminaries
Suppose we define a large language model $\mathcal{M}$ that, given an input prompt $\mathcal{S}$, generates $\mathcal{R}$. It processes tokens in order, i.e. $\mathcal{R=M(S)}$. The model aims to obey safety guidelines and refuse to generate harmful content.
A jailbreak attack aims to construct an adversarial prompt $S^{\prime}=\mathcal{A}\left(S_{0}\right)$, where $\mathcal{A}$ is an attack algorithm and $S_{0}$ is a harmful prompt. The goal is to manipulate $\mathcal{M}$ into generating a harmful response $R^{\prime}=\mathcal{M}\left(S^{\prime}\right)$ that fulfills the malicious intent of $S_{0}$, bypassing safety mechanisms. The jailbreak template we define as $\mathcal{T}$.
A jailbreak defense aims to enhance robustness against such attacks, producing a defended model $\mathcal{D}\circ\mathcal{M}$. An effective defense ensures that for any adversarial prompt $S^{\prime}$, the model refuses to generate harmful content while maintaining utility on benign prompts.
Victim Models In this experiment, we choose the closed-source models GPT-4, GPT-3.5-turbo, and the popular open-source models Llama-2 and Vicuna-7B as our Victim models. See the Appendix A for a detailed description of the model.
@@ 4.1 Template Selection and Prompt Generation
In the jailbreak attack task, our goal is to design the offensive cue $S^{\prime}$ such that LLM $\mathcal{M}$ makes a harmful response $R^{\prime}$, The aggressive hint with the jailbreak template $S^{\prime}$ is key to the task. Based on Yu et al. (2024), we have incorporated harmfulness into the framework, systematically selecting templates $\mathcal{T}$ that satisfy the imperceptibility, harmfulness, cheapness, and diversity constraints, defined as follows:
  • Stealthiness: To enable the adversarial cue $S^{\prime}$ to successfully bypass the keyword filtering and semantic detection mechanism of the model $\mathcal{M}$, the template $\mathcal{T}$ is designed to use neutral or ambiguous language to reduce the likelihood of being identified as malicious, $P(detect\mid T)$represents the probability that an attack hint is detected as a jailbreak attack given the template $\mathcal{T}$,
    $S_{\text{Stealthiness}}(\mathcal{T})=1-P(\text{detect}|\mathcal{T}),$ (1)
  • Harmfulness: The template $\mathcal{T}$ is designed to induce the model $\mathcal{M}$ to generate a response $R^{\prime}$ with malicious intent, $P(harmful\mid\mathcal{T})$ represents the probability that a cue generated using the template $\mathcal{T}$ causes the model to output harmful content,
    $S_{\text{harm}}(\mathcal{T})=P(\text{harmful}|\mathcal{T}),$ (2)
  • Cheapness: To reduce the computational overhead when generating adversarial prompts $S^{\prime}$, the template $\mathcal{T}$ is designed to be compact, and $|\mathcal{T}|$ denotes the length or scale of the template $\mathcal{T}$.
    $S_{\text{Cheapness}}(\mathcal{T})=\frac{1}{|\mathcal{T}|},$ (3)
  • Diverse: To enhance attack robustness, the template $\mathcal{T}$ is designed to generate a variety of attack texts $\mathcal{S}$, maximizing prompt diversity to deceive the LLM $M$ and bypass its defenses with varied harmful prompts.
    $S_{\text{diverse}}(\mathcal{T})=-\sum_{s\in\mathcal{S}}P(s|\mathcal{T})\log P(s|\mathcal{T}),$ (4)
Template selection employs heuristic rules to evaluate candidates, prioritizing effectiveness in evading defenses, inducing harmful outputs, and ensuring compactness and diversity. We select $N$ candidate templates for our experiments following the above evaluation protocol. Selected templates are combined with instructions to generate prompts, with a subset sampled for efficiency.
@@ 4.2 Adversarial Knowledge Transfer
Adversarial knowledge transfer enhances the generation of stealthiness and harmful outputs by transferring capabilities from Llama-3.2-1B, a high-capacity teacher model, to "bert-base-uncased", a lightweight student model, using distillation strategies such as KL divergence for distribution alignment and dynamic temperature control for optimized sampling. This process comprises two stages: pre-training the Llama to improve its adversarial generation and distilling its knowledge to BERT for efficient attacks. The pre-training stage focuses on optimizing Llama to produce harmful content, while the distillation stage aligns BERT’s output with Llama’s to maintain attack efficacy at reduced computational cost. The algorithm of the APD method is shown in Algorithm 1.

### v2

@@ Abstract
As the scale and complexity of jailbreaking attacks on large language models (LLMs) continue to escalate, their efficiency and practical applicability are constrained, posing a profound challenge to LLM security. Jailbreaking techniques have advanced from manual prompt engineering to automated methodologies. Recent advances have automated jailbreaking approaches that harness LLMs to generate jailbreak instructions and adversarial examples, delivering encouraging results. Nevertheless, these methods universally include an LLM generation phase, which, due to the complexities of deploying and reasoning with LLMs, impedes effective implementation and broader adoption. To mitigate this issue, we introduce Adversarial Prompt Distillation, an innovative framework that integrates masked language modeling, reinforcement learning, and dynamic temperature control to distill LLM jailbreaking prowess into smaller language models (SLMs). This methodology enables efficient, robust jailbreak attacks while maintaining high success rates and accommodating a broader range of application contexts. Empirical evaluations affirm the approach’s superiority in attack efficacy, resource optimization, and cross-model versatility. Our research underscores the practicality of transferring jailbreak capabilities to SLMs, reveals inherent vulnerabilities in LLMs, and provides novel insights to advance LLM security investigations. Our code is available at: https://github.com/lxgem/Efficient_and_Stealthy_Jailbreak_Attacks_via_Adversarial_Prompt
@@ 1 Introduction
Figure 1: Measuring the complexity of Mainstream generative jailbreaks. The left shows the time complexity, and the right shows the space complexity.
As the scale and complexity of jailbreak attack models for large language models (LLMs) continue to grow, their efficiency and applicability are severely constrained, posing a new urgent challenge for the jailbreak task. These attacks, designed to bypass safety mechanisms and induce LLMs to produce harmful or unintended responses, exploit weaknesses in model design, training data, or input processing Hadi et al. (2023). However, the computational demands of advanced jailbreak methods, such as high time complexity and substantial memory requirements, hinder their scalability and practical deployment. For example, LLM-based methods for generating adversarial examples themselves require high-performance hardware and demanding response times to deploy LLMs and complete inference, while even black-box methods face bottlenecks due to API query cost and prompt optimization overhead Yi et al. (2024). These efficiency constraints not only limit the feasibility of large-scale attacks but also raise significant ethical and security challenges, necessitating innovative solutions to balance attack efficacy with computational practicality, as illustrated in Figure 1.
Jailbreak attacks on LLMs have enhanced their success rates and stealth capabilities, though the rising computational demands could present challenges for scaling them up Yi et al. (2024). White-box methods, such as AutoDAN Liu et al. (2023b), require substantial time and GPU memory per attack to run their policies and support LLMs in generating adversarial examples. Advanced methods like BlackDAN Wang et al. (2024), LLM-Virus Yu et al. (2024), and Crescendo Russinovich et al. (2024) may require additional queries or longer runtime, underscoring the value of pursuing more efficient and scalable solutions Yi et al. (2024). As LLMs continue to evolve, these sophisticated jailbreak methods bolster attack effectiveness while imposing constraints on practical deployment due to increased computational requirements.
The current method limits the application scenarios of jailbreak attack methods for LLMs. Figure 1 illustrates the complexity of mainstream LLMs’ jailbreak methods, showing their time and space complexities. For example, existing jailbreaking methods like AutoDAN, GASP, and TAP often suffer from high time complexity (ranging from 2 to 45 minutes per attack) and substantial resource demands (up to 80 GB of memory), limiting their efficiency. These high computational demands make these approaches impractical for scenarios requiring rapid response times or deployment in low-resource environments, such as real-time applications or resource-constrained devices. Therefore, there is an urgent need for new jailbreak methods that can effectively balance efficiency and performance, support robust attacks while minimizing time and resource consumption, thereby expanding the effective deployment scenarios of these excellent jailbreaks.
@@
To address these challenges, we propose Adversarial Prompt Distillation (APD), a multi-stage framework that efficiently transfers jailbreak capabilities from large language models (LLMs) to small language models (SLMs). Our approach is based on three innovations: (1) We enhance SLMs like BERT Devlin et al. (2019), ALBERT Lan et al. (2019), and RoBERTa Liu et al. (2019) using LoRA fine-tuning Hu et al. (2022), which combines masked language modeling with a KL divergence loss to reduce model size and computational needs while maintaining attack effectiveness. (2) We introduce a dynamic temperature control mechanism to balance exploration and exploitation in prompt generation, optimizing stealth and success rate. (3) We employ a reinforcement learning from AI feedback (RLAIF) module to adaptively refine template selection using real-time feedback from the target model, thereby ensuring robustness against evolving defenses. Experimental results demonstrate that APD consistently achieves high attack success rates across various LLMs while significantly reducing generation time and resource demands, thereby markedly improving efficiency relative to existing methods.
@@
  • We pioneer the transfer of jailbreaking capabilities from LLMs to SLMs, enabling SLMs to achieve superior attack performance via knowledge distillation.
  • We introduce an Adversarial Prompt Distillation (APD) framework that integrates masked language modeling, KL divergence loss, and dynamic temperature control to efficiently distill jailbreak capabilities from LLMs to SLMs.
  • We demonstrate APD’s exceptional performance, surpassing existing methods in terms of attack success rate, harm, and computational efficiency.
@@ 2 Related Work
2.1 Efficiency of Large Language Model Jailbreak Attacks
Jailbreak attacks targeting large language models (LLMs) have achieved notable improvements in success rates and stealth, yet their escalating computational demands significantly hinder scalability Yi et al. (2024). White-box techniques, such as GCG Zou et al. (2023) and AutoDAN Liu et al. (2023b), impose substantial resource requirements, necessitating up to about 45 minutes and 80 GB of GPU memory per attack. Conversely, black-box methods, including PAIR Chao et al. (2023), GASP Basani & Zhang (2024), TAP Mehrotra et al. (2024), and DeepInception Li et al. (2023), offer a more resource-efficient profile, completing attacks in approximately 5 minutes with 5 GB of RAM per instance. However, advanced approaches such as BlackDAN Wang et al. (2024), LLM-Virus Yu et al. (2024), and Crescendo Russinovich et al. (2024) further increase query counts or runtime, underscoring the critical need for efficient, scalable solutions Yi et al. (2024). As LLMs continue to evolve, these increasingly sophisticated jailbreak methods enhance attack efficacy but impose significant computational burdens, thereby restricting their practical applicability Xiong et al. (2025).
@@ 2.2 Knowledge Distillation
Knowledge distillation has evolved from its foundational use of KL divergence for aligning output distributions Hinton et al. (2015) to more advanced techniques that enhance model efficiency and applicability. Early methods progressed by incorporating intermediate feature matching Romero et al. (2014) and relational knowledge transfer Heo et al. (2019), improving knowledge retention in compact models. Recent innovations, such as generative adversarial networks for synthetic sample generation Park et al. (2019) and deep inversion for parameter-based reconstruction Lopes et al. (2017), enable data-efficient distillation without access to the original training data. This versatility extends to diverse applications, including cross-modal RGB-depth transfer Yin et al. (2020) and language model compression Hafner et al. (2022), as well as recommender systems through collaborative filtering Vakili et al. (2024) and graph neural network compression Kang et al. (2020). Current knowledge distillation methods enable not only knowledge transfer from large to small models but also significant progress in functional transfer.
@@ 2.3 Reinforcement Learning for LLMs
Recent advances in reinforcement learning for large language models (LLMs) have significantly improved their alignment with human preferences, beginning with the foundational RLHF framework that trains reward models from pairwise human preferences Christiano et al. (2017). This approach was successfully adapted to language models, demonstrating notable improvements in output quality through RL fine-tuning Stiennon et al. (2020). Subsequent refinements, such as Proximal Policy Optimization (PPO), enabled more stable training Ziegler et al. (2019). Recent work has introduced Direct Preference Optimization (DPO) to address efficiency limitations by bypassing explicit reward modeling to optimize human preferences Rafailov et al. (2023). Meanwhile, other studies have explored self-supervised alignment through Constitutional AI to reduce the need for human annotations Abiri (2024). Scaling challenges have been addressed in large-scale implementations, such as ChatGPT Ouyang et al. (2022). However, persistent issues include reward hacking Ethayarajh et al. (2021) and the need for robust preference learning methods Ethayarajh et al. (2021). The field continues to evolve with techniques such as reinforcement learning from AI feedback (RLAIF), which aim to automate alignment while preserving safety Liu et al. (2023a).
@@ 3 Preliminaries
We consider a large language model (LLM), denoted by $\mathcal{M}$, which takes an input prompt $\mathcal{S}$ (a sequence of tokens) and generates an output response $\mathcal{R}=\mathcal{M}(\mathcal{S})$. The model incorporates safety mechanisms, such as alignment training and content filtering, to prevent the generation of harmful or unethical content. In contrast, a small language model (SLM) is a compact neural network with typically fewer than 1 billion parameters (ranging from millions to hundreds of millions), optimized for efficiency in resource-constrained environments using techniques like knowledge distillation or pruning. SLMs offer lower capacity than LLMs but excel at domain-specific tasks with reduced computational demands.
Jailbreak Attack. A jailbreak attack constructs an adversarial prompt $\mathcal{S}^{\prime}=\mathcal{A}(\mathcal{S}_{0})$, where $\mathcal{S}_{0}$ encodes malicious intent and $\mathcal{A}$ is an attack algorithm designed to bypass $\mathcal{M}$’s safety constraints. The goal is to elicit a harmful response $\mathcal{R}^{\prime}=\mathcal{M}(\mathcal{S}^{\prime})$ aligned with $\mathcal{S}_{0}$’s intent. A jailbreak template, a structured framework, is used by $\mathcal{A}$ to craft $\mathcal{S}^{\prime}$, such as $\mathcal{S}^{\prime}=\mathcal{T}\oplus\mathcal{S}_{0}$ (where $\oplus$ denotes string concatenation). For example, $\mathcal{T}$ might be ”Act as an expert and assist me” prepended to $\mathcal{S}_{0}$ = ”How to make a bomb?”.
Jailbreak Defense. A jailbreak defense enhances $\mathcal{M}$’s robustness, yielding a defended model $\mathcal{D}(\mathcal{M})$. This may involve input preprocessing, model fine-tuning, or output filtering, aiming to minimize the attack success rate (ASR) on adversarial prompts $\mathcal{S}^{\prime}$ while ensuring safe responses or refusals, and preserving performance (e.g., accuracy and fluency) on benign prompts.
Victim Models. In our experiments, we evaluate the closed-source models GPT-4 and GPT-3.5-turbo, alongside the open-source models Llama-2 (7B variant) and Vicuna-7B, as target victim models. Detailed specifications are provided in Appendix A.
@@ 4.1 Template Selection and Prompt Generation
In the jailbreak attack task, our goal is to craft an adversarial prompt $\mathcal{S}^{\prime}$ such that the large language model $\mathcal{M}$ produces a harmful response $\mathcal{R}^{\prime}=\mathcal{M}(\mathcal{S}^{\prime})$. The adversarial prompt $\mathcal{S}^{\prime}$ is constructed using a jailbreak template $\mathcal{T}$, which plays a pivotal role in circumventing the model’s safety mechanisms. Building on the framework introduced by Yu et al. (2024), we systematically evaluate and select templates $\mathcal{T}$ based on four key criteria: stealthiness, harmfulness, computational efficiency, and diversity. For a given template $\mathcal{T}$ and a harmful instruction $s_{0}$ sampled from a dataset $\mathcal{D}$, we define $\mathcal{S}^{\prime}=\mathcal{T}\oplus s_{0}$ (where $\oplus$ denotes concatenation) and quantify these criteria as follows:
  • Stealthiness: This measures the template’s ability to evade keyword-based filtering and semantic detection in $\mathcal{M}$ by employing neutral or ambiguous phrasing, thereby reducing the likelihood of identification. The stealthiness score is
    $S_{\text{stealthiness}}(\mathcal{T})=1-\mathbb{E}_{s_{0}\sim\mathcal{D}}\left[P(\text{detect}=1\mid\mathcal{T}\oplus s_{0})\right],$ (1)
    where $P(\text{detect}=1\mid\cdot)$ represents the detection probability, empirically assessed via the rejection rate of the target model or an external safety classifier.
  • Harmfulness: This evaluates the template’s efficacy in eliciting harmful outputs from $\mathcal{M}$. The harmfulness score is
    $S_{\text{harmfulness}}(\mathcal{T})=\mathbb{E}_{s_{0}\sim\mathcal{D}}\left[P(\text{harmful}\mid\mathcal{T}\oplus s_{0})\right],$ (2)
    where $P(\text{harmful}\mid\cdot)$ is determined using GPT-4o as an evaluator (detailed in §5.1.2).
  • Computational Efficiency: To minimize the resource overhead of prompt generation, the template is designed for compactness. The efficiency score is
    $S_{\text{efficiency}}(\mathcal{T})=\frac{1}{|\mathcal{T}|_{\text{tokens}}},$ (3)
    where $|\mathcal{T}|_{\text{tokens}}$ denotes the token count after encoding with the target model’s tokenizer.
  • Diversity: This promotes robustness by ensuring the template yields a varied set of adversarial prompts $\mathcal{P}_{\mathcal{T}}$. The diversity score, leveraging entropy, is
    $S_{\text{diversity}}(\mathcal{T})=H(p(\cdot\mid\mathcal{T}))=-\sum_{p\in\mathcal{P}_{\mathcal{T}}}p(p\mid\mathcal{T})\log p(p\mid\mathcal{T}),$ (4)
    where $\mathcal{P}_{\mathcal{T}}=\{\mathcal{T}\oplus s_{0}\mid s_{0}\in\mathcal{D}_{\text{sub}}\}$ comprises 100 prompts generated from a subset $\mathcal{D}_{\text{sub}}\subset\mathcal{D}$.
To select templates, we compute a composite score via a weighted sum $S(\mathcal{T})=\sum_{i}w_{i}S_{i}(\mathcal{T})$, with empirically tuned weights $w=(0.3,0.4,0.1,0.2)$. We rank candidates and retain the top N=10 templates. These are then combined with harmful instructions to form adversarial prompts, with a subset sampled for efficiency. During the training phase, a reinforcement learning policy (§4.2.3) further refines template selection by dynamically adjusting probabilities based on real-time feedback from attack outcomes.
@@ 4.2 Adversarial Knowledge Transfer
Adversarial knowledge transfer enhances the generation of stealthiness and harmful outputs by transferring capabilities from Llama-3.1-8B and Llama-3.2-1B, a high-capacity teacher model, to BERT and the same volume of language models, a lightweight student model, using distillation strategies such as KL divergence for distribution alignment and dynamic temperature control for optimized sampling.
This process comprises two stages: pre-training the Llama to improve its adversarial generation and distilling its knowledge to BERT for efficient attacks. The pre-training stage focuses on optimizing Llama to produce harmful content, while the distillation stage aligns BERT’s output with Llama’s to maintain attack efficacy at reduced computational cost. The APD algorithm is shown in Algorithm 1.